OpenStudy (lgbasallote):

an object moves along the x axis according to the equation x = 3.00t^2 - 2.00t + 3.00 where x is in meters and t is in seconds. Determine the average speed between t = 2.00 s and t = 3.00s do i just substitute t again? @mathslover :) just help if you can hehe

mathslover (mathslover):

@lgbasallote The average speed is the distance passed divided by the time it took : At T=2 the object was at 3*2^2-2*2+3 = 11m . At T=3 the object was 3*3^2-2*3+3 = 24m. Total distance passed is 24-11 = 13 meters. It took one second . So the average speed is 13/1 = 13m/s.

mathslover (mathslover):

that is the answer...

OpenStudy (lgbasallote):

distance passed is just substituting t right?

OpenStudy (lgbasallote):

what about for instantaneous acceleration??

mathslover (mathslover):

The acceleration function is the derivative of the speed function(which we found in B). So : a(t) = v'(t) = 6. this function is constant , so the instantaneous acceleration is always 6m/s^2 at any time.

OpenStudy (lgbasallote):

uhh is there a non-calculus way??

OpenStudy (mimi_x3):

the only way to find the acceleration is differentiating the velocity.

OpenStudy (lgbasallote):

wait..i skipped a letter...what is the instantaneous speed at t = 2 and t = 3? is it \[\frac{3(2)^2 - 2(2) + 3}{2}\] for t=2?

OpenStudy (lgbasallote):

and acceleration = \(\frac{d^2r}{dt^2} (3t^2 - 2t + 3)\) right?

OpenStudy (lgbasallote):

i mean you got the acceleration by taking the double derivative of 3t^2 - 2t + 3?

OpenStudy (mimi_x3):

Yes, the when differentiating the displacement you get the velocity then you differentiate the velocity to get the acceleration.

OpenStudy (lgbasallote):

@mathslover did i do my instantaneous speed right?

mathslover (mathslover):

Instantaneous speed at time t is the derivative of the location function at the point t. so we need the derive the function : x(t) = 3*t^2-2*t+3 so v(t) = x'(t) = 6t-2 . for the time t=2 we get 6*2-2 = 10m/s . for the time t=3 we get 6*3-2 = 16m/s Note that the answer for A is actually the average of the two answers of B . That is because the acceleration here is constant but that is not always the case! That is why the form to solve A is more general and will always give you the correct result even when this is not the case.

OpenStudy (lgbasallote):

do i ALWAYS use derivatives??

mathslover (mathslover):

physics can include derivative in any question

mathslover (mathslover):

that is why CALCULUS is the must part of maths and physics if no maths than no physics :(

OpenStudy (lgbasallote):

hmm instantaneous acceleration for t = 2 and t = 3 are both 6 right? since a is constant

mathslover (mathslover):

very great .... congr8s u r going right

OpenStudy (lgbasallote):

last follow up question is...what time is the object at rest? i assume i equate 3t^2 - 2t + 3 = 0 then solve for t?

OpenStudy (lgbasallote):

oh wait!!! 6t - 2 = 0!!

OpenStudy (lgbasallote):

because at rest means velocity is 0

mathslover (mathslover):

right...

OpenStudy (lgbasallote):

so when t = 1/3 seconds the object is at rest?

mathslover (mathslover):

right !!

OpenStudy (lgbasallote):

yay thanks!

mathslover (mathslover):

:) my pleasure.... yw:) ....mention not ....and thanks for tagging me here :)